OpenStudy (anonymous):

The line y = 2x + 3 is tangent to the parabola y = x2 + B. Find B. Can someone explain this step by step

OpenStudy (freckles):

so y=x^2+B => y'=2x+0 =>y'=2x The slope of the tangent line is 2 so 2x=2 => x=1 =>y=2(1)+3=5 =>5=1^2+B =>5-1=B =>B=4

OpenStudy (anonymous):

ahhh makes sense now

OpenStudy (freckles):

awesome! :)

OpenStudy (anonymous):

I just dont get why you plug that x back into the tangent line equation then set the it equal to the b equation

OpenStudy (anonymous):

I'm trying to the connect the relationship but idk

OpenStudy (freckles):

ok well i know that (x,y) is on both y=x^2+B and y=2x+3

OpenStudy (freckles):

well the point that they touch I'm calling (x,y)

OpenStudy (freckles):

we knew x

OpenStudy (freckles):

from setting the derivative=slope of the tangent line and solving for x

OpenStudy (freckles):

so i had enough info to find out what y was on y=2x+3

OpenStudy (freckles):

which would also be the same y on y=x^2+B

OpenStudy (anonymous):

AH thats right the derivative of that equation is the slope at that point?

OpenStudy (freckles):

Yep! :)

OpenStudy (anonymous):

SWEET thank you! I wish I could give you two medals for this lol!

OpenStudy (accessdenied):

i give medal too -- theres your second medal! :D

OpenStudy (freckles):

I don't need medals. I just need your love (platonic love of course). :)

OpenStudy (freckles):

Thanks you guys. You are so sweet.

OpenStudy (anonymous):

You has my math thats more than my love.